Please help me with this, I don't see in my book where they discused how to do this.
"A machinist drilled a cone-shaped hole into a solid cube of metal as shown. If the cube's sides have a length of 4 centimeters, what is the volume of the metal cube after the cone is drilled? Use 3.14 for pi and round your answer to the nearest tenth."
Help would be MUCH appreciated!! Thank you!!!!!! :)

Answer by stanbon(57424)

"A machinist drilled a cone-shaped hole into a solid cube of metal as shown. If the cube's sides have a length of 4 centimeters, what is the volume of the metal cube after the cone is drilled? Use 3.14 for pi and round your answer to the nearest tenth."
-------------------
Cube Volume = 4^3 = 64 cm^3
----
Cone Volume = (1/3)(3.14)(radius^2)*4
----
You did not post the radius of the cone.
-------------------------------------------------------------------------
Put that in the "Cone Volume" formula to get the volume of the cone.
---
Then subtract the volume of the cone from the volume of the cube
to get the answer you want.
==============================
cheers,
Stan H.